The average lifetime of mu-mesons with a speed of
Answer #1

6) we use the following equation

    \Delta t=\gamma \Delta t_{o}---(1)

where \Delta t is the average time in the rest frame, \Delta t_{o} is the proper time for muon moving with velocity 0.95c

\Rightarrow \Delta t=\frac{\Delta t_{0}}{\sqrt{1-\frac{v^{2}}{c^{2}}}}

\Rightarrow \Delta t=\frac{6*10^{-6}}{\sqrt{1-{0.95^{2}}}}=1.92*10^{-5}sec

7) Classically:

    d=v*t=(0.8c)*(1.8*10^{-8})=4.32m

   Relativistically:

   d=v*t'=v*\gamma t=v*\frac{t}{\sqrt{1-\frac{v^{2}}{c^{2}}}}=0.8c*\frac{1.8*10^{-8}}{\sqrt{1-(0.8)^{2}}}=7.2m

8) the velocity of A relative to the earth (u)=0.8c

    the velocity of B relative to earth (v)= -0.6c

    the velocity of A measured from B is

   {u}'=\frac{u-v}{1-\frac{uv}{c^{2}}}=\frac{0.8c-(-0.6c)}{1-\frac{0.8c*(-0.6c)}{c^{2}}}=0.96c

9) the speed of nucleus relative to lab (v)=0.5c

   the speed of electron relative to nucleus ({u}')=0.8c

   the velocity of electron in lab frame (u) is given by

u=\frac{{u}'+v}{1+\frac{{u}'v}{c^{2}}}=\frac{0.8c+0.5c}{1+(0.8*0.5)}=0.93c

10) the observed wavelength is \lambda _{obs}=589+10=599nm

      \lambda _{source}=589nm

    we use the doppler relation

     \lambda _{obs}=\lambda _{source}\sqrt{\frac{1+\frac{v}{c}}{1-\frac{v}{c}}}

\Rightarrow \left [ \frac{\lambda _{obs}}{\lambda _{sou}} \right ]^{2}={\frac{1+\frac{v}{c}}{1-\frac{v}{c}}}

\Rightarrow \left [ \frac{599}{589} \right ]^{2}={\frac{1+\frac{v}{c}}{1-\frac{v}{c}}}

Simplyfying the above equation we get

\frac{v}{c}=0.017